# Learning with Local Search MCMC Layers

### v1

@@ Appendix B Additional material
B.1 Fenchel-Young loss for $K=1$ in the unsupervised setting
This proposition is analogous to Section 4.3, but in the unsupervised setting, when using a data-based initialization method – i.e., the original CD initialization scheme, without persistent Markov chains. See Section B.2 for a detailed discussion about this. {proposition} Let $\bm{p}^{(1)}_{\bm{\theta},\bar{Y}_{N}}$ denote the distribution of the first iterate of the Markov chain defined by the Markov transition kernel given in Eq. 3, with proposal distribution $q$ and initialized by ${\bm{y}}^{(0)}={\bm{y}}_{i}$, with $i\sim\mathcal{U}(\llbracket 1,N\rrbracket)$. There exists a dataset-dependent regularization $\Omega_{\bar{Y}_{N}}$ with the following properties: $\Omega_{\bar{Y}_{N}}$ is $tN/\sum_{i=1}^{N}\mathbb{E}_{q({\bm{y}}_{i},\,\cdot\,)}||Y-{\bm{y}}_{i}||_{2}^{2}$-strongly convex; it is such that:
@@ B.1 Fenchel-Young loss for $K=1$ in the unsupervised setting
and the Fenchel-Young loss $L_{\Omega_{\bar{Y}_{N}}}$ generated by $\Omega_{\bar{Y}_{N}}$ is $\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}_{q({\bm{y}}_{i},\,\cdot\,)}||Y-{\bm{y}}_{i}||_{2}^{2}/t$-smooth in its first argument, and such that $\nabla_{\bm{\theta}}L_{\Omega_{\bar{Y}_{N}}}(\bm{\theta}\,;{\bm{y}})=\mathbb{E}_{\bm{p}^{(1)}_{\bm{\theta},\bar{Y}_{N}}}\left[Y\right]-{\bm{y}}$.
The proof is given in Section C.7.
B.2 Markov chain initialization
In contrastive divergence (CD) learning, the intractable expectation in the log-likelihood gradient is approximated by short-run MCMC, initialized at the data distribution [22] (using a Gibbs sampler in the setting of Restricted Boltzmann Machines).
@@ B.2 Markov chain initialization
[…]
@@
[…]
However, neither of these can be applied to the supervised (or conditional) setting, as observed in [38] in the context of conditional Restricted Boltzmann Machines (which are a type of EBMs). Indeed, on the one hand, PCD takes advantage of the fact that the parameter $\hat{\bm{\theta}}$ does not change too much from one iteration to the next, so that a Markov Chain that has reached equilibrium on $\hat{\bm{\theta}}_{n}$ is not far from equilibrium on $\hat{\bm{\theta}}_{n+1}$. This does not hold in the supervised setting, as each $\bm{x}_{i}$ leads to a different $\hat{\bm{\theta}}_{i}=g_{\hat{W}}(\bm{x}_{i})$. On the other hand, the data-based initialization method in CD would amount to initialize the chains at the empirical marginal data distribution on $\mathcal{Y}$, and would be irrelevant in a supervised setting, since the distribution we want each Markov Chain to approximate is conditioned on the input $\bm{x}_{i}$.
An option is to use persistent chains if training for multiple epochs, and to initialize the Markov Chain associated to $(\bm{x}_{i},{\bm{y}}_{i})$ for epoch $j$ at the final state of the one associated to the same data point $(\bm{x}_{i},{\bm{y}}_{i})$ at epoch $j-1$. However, this method is relevant than PCD in the unsupervised setting, as $\hat{w}$ changes a lot more in a full epoch than $\hat{\theta}$ in just one gradient step in the unsupervised setting. It might be relevant, however, if each epoch consists in a single, full-batch gradient step. Nevertheless, it would require to store a significant number of states $\bm{y}_{i}^{(n,\,K)}$ (one for each point in the dataset). The solution we propose, for both full-batch and mini-batch settings, is to initialize the chains at the empirical data distribution conditioned on the input $\bm{x}_{i}$, which amounts to initialize them at the ground-truth ${\bm{y}}_{i}$.
This discussion is summed up in Table 4.
@@
Init. [-0.2ex] Method Setting Unsupervised Supervised, Batch Supervised, Mini-Batch
[…]
@@
B.3 Proposal distribution design for the DVRPTW
@@
The selected moves, designed for Local Search algorithms on vehicle routing problems (specifically for the PC-VRPTW for serve request and remove request), are given in Table 5.
@@
In order to transform these deterministic moves into proposals, we first adapt the choice of clients $i$ and $j$, by sampling $i$ uniformly from $V^{1}_{s}(\bm{y})$, which contains the set of valid choices of client $i$ for move $s$ from solution $\bm{y}$. Then, we sample $j$ from $V^{2}_{s}(\bm{y})[i]\setminus\{i\}$ using the following softmax distribution: $P_{s}(j\mid i)=\frac{\exp\left[-d(i,j)/\beta\right]}{\sum_{k\in V^{2}_{s}(\bm{y})[i]\setminus\{i\}}\exp\left[-d(i,k)/\beta\right]}$, where $\beta>0$ is a neighborhood sampling temperature. The set $V^{2}_{s}(\bm{y})[i]$ contains all valid choices of client $j$ for move $s$ from solution $\bm{y}$, and is precised along with $V^{1}_{s}(\bm{y})$ in Table 6. We normalize the distance measures inside the softmax, by dividing them by the maximum distance: $d(i,\cdot)\leftarrow d(i,\cdot)/\max_{k\in V^{2}_{s}(\bm{y})[i]\setminus\{i\}}d(i,k)$.
@@
relocate pair $\mathcal{D}(\bm{y})\setminus\bigl{\{}\mathcal{D}_{2}(\bm{y})\cup\mathcal{D}^{\text{last}}(\bm{y})\bigr{\}}$ $\mathcal{D}(\bm{y})\setminus\bigl{\{}\mathcal{D}_{2}(\bm{y})\cup\mathcal{D}^{\text{last}}(\bm{y})\cup\{\mathsf{next}(i)\}\bigr{\}}$
@@
serve request $\overline{\mathcal{D}}(\bm{y})$ $\mathcal{D}(\bm{y})\cup\mathcal{I}_{D}(\bm{y})$
@@
Table 6: Sets of valid clients for each move. $\mathcal{D}(\bm{y})$ contains all dispatched clients in solution $\bm{y}$. $\mathcal{D}_{1}(\bm{y})$ contains all dispatched clients that are the only client in their route. $\mathcal{D}_{2}(\bm{y})$ contains all dispatched clients that are in a route with $2$ clients or less. $\mathcal{D}^{\text{last}}(\bm{y})$ contains all dispatched clients that are the last of their route. $\overline{\mathcal{D}}(\bm{y})$ contains all non-dispatched clients. $\mathcal{I}_{D}(\bm{y})$ contains the depot of the first empty route, if it exists (all routes may be non-empty), or else is the empty set. $\mathcal{I}_{1}(\bm{y})$ contains the only client in the last non-empty route if it contains exactly one client, or else is the empty set.
@@
  • For swap pair, the same arguments hold (leading to the same form for $q_{s}$), except for the equalities $V^{2}_{s}(\bm{y})[i]=V^{2}_{s}(\bm{y}^{\prime})[i]$ and $V^{2}_{s}(\bm{y})[j]=V^{2}_{s}(\bm{y}^{\prime})[j]$. Thus, the ratio $\frac{q_{s}(\bm{y}^{\prime},\bm{y})}{q_{s}(\bm{y},\bm{y}^{\prime})}$ must be computed explicitly, which has complexity $\mathcal{O}(\mathcal{R}^{\omega})$.
@@
    $q_{s}(\bm{y},\bm{y}^{\prime})$ =|$=\frac{|\bigl{\{}\mathcal{D}(\bm{y})\setminus\mathcal{D}_{1}(\bm{y})\bigr{\}}\cup\mathcal{I}_{1}(\bm{y})|}{|\bigl{\{}\mathcal{D}(\bm{y})\setminus\mathcal{D}_{1}(\bm{y})\bigr{\}}\cup\mathcal{I}_{1}(\bm{y})|+|\overline{\mathcal{D}}(\bm{y})|}\times\frac{1}{|\bigl{\{}\mathcal{D}(\bm{y})\setminus\mathcal{D}_{1}(\bm{y})\bigr{\}}\cup\mathcal{I}_{1}(\bm{y})|}$
    $=\frac{1}{|\bigl{\{}\mathcal{D}(\bm{y})\setminus\mathcal{D}_{1}(\bm{y})\bigr{\}}\cup\mathcal{I}_{1}(\bm{y})|+|\overline{\mathcal{D}}(\bm{y})|}$
@@
    $q_{s}(\bm{y}^{\prime},\bm{y})$ $=\frac{|\overline{\mathcal{D}}(\bm{y}^{\prime})|}{|\bigl{\{}\mathcal{D}(\bm{y}^{\prime})\setminus\mathcal{D}_{1}(\bm{y}^{\prime})\bigr{\}}\cup\mathcal{I}_{1}(\bm{y}^{\prime})|+|\overline{\mathcal{D}}(\bm{y}^{\prime})|}$
    ×$\times\frac{\exp\left[-\bar{d}(i)/\beta\right]}{\exp\left[-\bar{d}(i)/\beta\right]+\sum_{k\in\mathcal{D}(\bm{y}^{\prime})}\exp\left[-d(i,k)/\beta\right]}$
    $=\frac{|\overline{\mathcal{D}}(\bm{y})|+1}{|\bigl{\{}\mathcal{D}(\bm{y})\setminus\mathcal{D}_{1}(\bm{y})\bigr{\}}\cup\mathcal{I}_{1}(\bm{y})|+|\overline{\mathcal{D}}(\bm{y})|}$
    ×$\times\frac{\exp\left[-\bar{d}(i)/\beta\right]}{\exp\left[-\bar{d}(i)/\beta\right]+\sum_{\begin{subarray}{c}k\in\mathcal{D}(\bm{y})\\ k\neq i\end{subarray}}\exp\left[-d(i,k)/\beta\right]}.$
@@
    $q_{s}(\bm{y}^{\prime},\bm{y})$ $=\frac{|\overline{\mathcal{D}}(\bm{y}^{\prime})|}{|\bigl{\{}\mathcal{D}(\bm{y}^{\prime})\setminus\mathcal{D}_{1}(\bm{y}^{\prime})\bigr{\}}\cup\mathcal{I}_{1}(\bm{y}^{\prime})|+|\overline{\mathcal{D}}(\bm{y}^{\prime})|}$
    ×$\times\frac{\frac{1}{2}\cdot\exp\left[-d(i,\mathsf{prev}(i))\right]+\frac{1}{2}\cdot\exp\left[-d(i,\mathsf{next}(i))\right]}{\mathbf{1}_{\{\mathcal{I}_{D}(\bm{y}^{\prime})\neq\emptyset\}}\cdot\exp\left[-\bar{d}(i)/\beta\right]+\sum_{k\in\mathcal{D}(\bm{y}^{\prime})}\exp\left[-d(i,k)/\beta\right]}$
    $=\frac{|\overline{\mathcal{D}}(\bm{y})|+1}{|\bigl{\{}\mathcal{D}(\bm{y})\setminus\mathcal{D}_{1}(\bm{y})\bigr{\}}\cup\mathcal{I}_{1}(\bm{y})|+|\overline{\mathcal{D}}(\bm{y})|}$
    ×$\times\frac{\frac{1}{2}\cdot\exp\left[-d(i,\mathsf{prev}(i))\right]+\frac{1}{2}\cdot\exp\left[-d(i,\mathsf{next}(i))\right]}{\mathbf{1}_{\{\mathcal{I}_{D}(\bm{y}^{\prime})\neq\emptyset\}}\cdot\exp\left[-\bar{d}(i)/\beta\right]+\sum_{\begin{subarray}{c}k\in\mathcal{D}(\bm{y})\\ k\neq i\end{subarray}}\exp\left[-d(i,k)/\beta\right]}.$
@@
    $q_{s}(\bm{y},\bm{y}^{\prime})$ $=\frac{|\overline{\mathcal{D}}(\bm{y})|}{|\bigl{\{}\mathcal{D}(\bm{y})\setminus\mathcal{D}_{1}(\bm{y})\bigr{\}}\cup\mathcal{I}_{1}(\bm{y})|+|\overline{\mathcal{D}}(\bm{y})|}$
    ×$\times\frac{\frac{1}{2}\cdot\exp\left[-d(i,j)\right]+\frac{1}{2}\cdot\exp\left[-d(i,j^{\prime})\right]}{\mathbf{1}_{\{\mathcal{I}_{D}(\bm{y})\neq\emptyset\}}\cdot\exp\left[-\bar{d}(i)/\beta\right]+\sum_{k\in\mathcal{D}(\bm{y})}\exp\left[-d(i,k)/\beta\right]}$
@@
    $q_{s}(\bm{y},\bm{y}^{\prime})$ $=\frac{|\overline{\mathcal{D}}(\bm{y})|}{|\bigl{\{}\mathcal{D}(\bm{y})\setminus\mathcal{D}_{1}(\bm{y})\bigr{\}}\cup\mathcal{I}_{1}(\bm{y})|+|\overline{\mathcal{D}}(\bm{y})|}$
    ×$\times\frac{\exp\left[-\bar{d}(i)/\beta\right]}{\exp\left[-\bar{d}(i)/\beta\right]+\sum_{k\in\mathcal{D}(\bm{y})}\exp\left[-d(i,k)/\beta\right]}$
@@
    $q_{s}(\bm{y}^{\prime},\bm{y})=\frac{1}{|\bigl{\{}\mathcal{D}(\bm{y})\setminus\mathcal{D}_{1}(\bm{y})\bigr{\}}\cup\mathcal{I}_{1}(\bm{y})|+|\overline{\mathcal{D}}(\bm{y})|}.$
@@
In each case, we set $d(i,D)=+\infty$ to account for the fact that the depot can never be sampled during the process (except in the serve request / remove request move, where we allow the depot of the first empty route / last non-empty route to be selected, for which we use the average distance to other requests as explained earlier) – in fact, the distance measure from a client to the depot is not even defined in the original HGS implementation.

### v2

@@ Appendix B Additional material
B.1 Fenchel-Young loss for $K=1$ in the unconditional setting
[…]
@@ B.1 Fenchel-Young loss for $K=1$ in the unconditional setting
[…]
The proof is given in Section˜D.7.
B.2 Properties of the expected first iterate
{proposition}
Let ${\bm{\theta}}\in{\mathbb{R}}^{d}$, ${\bm{y}}\in{\mathcal{Y}}$. Let
${\mathcal{N}}_{\text{better}}({\bm{y}})\coloneqq\left\{{\bm{y}}^{\prime}\in{\mathcal{N}}({\bm{y}})\mid\langle{\bm{\theta}},{\bm{y}}^{\prime}\rangle+\varphi({\bm{y}}^{\prime})>\langle{\bm{\theta}},{\bm{y}}\rangle+\varphi({\bm{y}})\right\}$
denote the set of improving neighbors of $\bm{y}$} for the unregularized objective function. We have the following properties:
${\mathbb{E}}_{\bm{p}^{(1)}_{\bm{\theta},{\bm{y}}}}[Y]$ $\in\operatorname{\mathsf{conv}}\left(\mathcal{N}({\bm{y}})\cup\{{\bm{y}}\}\right),$
${\mathbb{E}}_{\bm{p}^{(1)}_{\bm{\theta},{\bm{y}}}}[Y]$ $\xrightarrow[t\rightarrow 0^{+}]{}{\bm{y}}+\!\!\!\!\!\sum_{{\bm{y}}^{\prime}\in{\mathcal{N}}_{\text{better}}({\bm{y}})}q({\bm{y}},{\bm{y}}^{\prime})\cdot({\bm{y}}^{\prime}-{\bm{y}}),$
$\text{and}\quad{\mathbb{E}}_{\bm{p}^{(1)}_{\bm{\theta},{\bm{y}}}}[Y]$ $\xrightarrow[t\rightarrow\infty]{}{\bm{y}}+\!\!\!\sum_{{\bm{y}}^{\prime}\in{\mathcal{N}}({\bm{y}})}\min\left[q({\bm{y}},{\bm{y}}^{\prime}),q({\bm{y}}^{\prime},{\bm{y}})\right]\cdot({\bm{y}}^{\prime}-{\bm{y}}).$
The proof is given in Section˜D.8. Thus, as the set ${\mathcal{N}}_{\text{better}}$ is defined according the value of the original, unregularized objective function ${\bm{y}}\mapsto\langle{\bm{\theta}},{\bm{y}}\rangle+\varphi({\bm{y}})$, the low temperature behavior of the regularized maximizer ${\mathbb{E}}_{\bm{p}^{(1)}_{\bm{\theta},{\bm{y}}}}[Y]$ effectively reflects the fact that the regularization function $\Omega_{\bm{y}}$ extends the influence of $\varphi$ from the vertices ${\mathcal{N}}({\bm{y}})\cup\{{\bm{y}}\}$ to their convex hull.
B.3 Markov chain initialization
In contrastive divergence (CD) learning, the intractable expectation in the log-likelihood gradient is approximated by short-run MCMC, initialized at the data distribution [Hinton, 2000] (using a Gibbs sampler in the setting of Restricted Boltzmann Machines).
@@ B.3 Markov chain initialization
for theconditional setting, with $B_{n}$ being the mini-batch (or full batch) used at iteration $n$, ${\bm{y}}_{i}$ the ground-truth structure associated to $\bm{x}_{i}$ in the dataset, and $\bm{y}_{i}^{(n+1,\,k)}$ the $k$-th iterate of Algorithm˜1, with maximization direction $g_{\hat{W}_{n}}(\bm{x}_{i})$, and initialization point $\bm{y}_{i}^{(n+1,\,0)}$. We also note:
@@
for the unconditional setting, with $\bm{y}^{(n+1,\,k)}$ being the $k$-th iterate of Algorithm˜1, with maximization direction $\hat{\bm{\theta}}_{n}$, and initialization point $\bm{y}^{(n+1,\,0)}$.
In CD learning of unconditional EBMs (i.e., in our unconditional setting), the Markov Chain is initialized at the empirical data distribution [Hinton, 2000, Carreira-Perpiñán and Hinton, 2005], as explained earlier. Persistent Contrastive Divergence (PCD) learning [Tieleman, 2008] modifies CD by maintaining a persistent Markov chain. Thus, instead of initializing the chain from the data distribution in each iteration, the chain continues from its last state in the previous iteration, by setting $\bm{y}^{(n+1,\,0)}=\bm{y}^{(n,\,K)}$. This approach aims to provide a better approximation of the model distribution and to reduce the bias introduced by the initialization of the Markov chain in CD. These are two types of informative initialization methods, which aim at reducing the mixing times of the Markov Chains.
However, neither of these can be applied to the conditional setting, as observed in [Mnih et al., 2012] in the context of conditional Restricted Boltzmann Machines (which are a type of EBMs). Indeed, on the one hand, PCD takes advantage of the fact that the parameter $\hat{\bm{\theta}}$ does not change too much from one iteration to the next, so that a Markov Chain that has reached equilibrium on $\hat{\bm{\theta}}_{n}$ is not far from equilibrium on $\hat{\bm{\theta}}_{n+1}$. This does not hold in the conditional setting, as each $\bm{x}_{i}$ leads to a different $\hat{\bm{\theta}}_{i}=g_{\hat{W}}(\bm{x}_{i})$. On the other hand, the data-based initialization method in CD would amount to initialize the chains at the empirical marginal data distribution on $\mathcal{Y}$, and would be irrelevant in a conditional setting, since the distribution we want each Markov Chain to approximate is conditioned on the input $\bm{x}_{i}$.
An option is to use persistent chains if training for multiple epochs, and to initialize the Markov Chain associated to $(\bm{x}_{i},{\bm{y}}_{i})$ for epoch $j$ at the final state of the one associated to the same data point $(\bm{x}_{i},{\bm{y}}_{i})$ at epoch j1j-1. However, this method is relevant than PCD in the unconditional setting, as $\hat{w}$ changes a lot more in a full epoch than $\hat{\theta}$ in just one gradient step in the unconditional setting. It might be relevant, however, if each epoch consists in a single, full-batch gradient step. Nevertheless, it would require to store a significant number of states $\bm{y}_{i}^{(n,\,K)}$ (one for each point in the dataset). The solution we propose, for both full-batch and mini-batch settings, is to initialize the chains at the empirical data distribution conditioned on the input $\bm{x}_{i}$, which amounts to initialize them at the ground-truth ${\bm{y}}_{i}$.
This discussion is summed up in Table˜4.
@@
Init. [-0.2ex] Method Setting Unconditional Conditional, Batch Conditional, Mini-Batch
Persistent $\bm{y}^{(n+1,0)}=\bm{y}^{(n,K)}$ $\bm{y}_{i}^{(n+1,\,0)}=\bm{y}_{i}^{(n,\,K)}$ /
Data-Based $\bm{y}^{(n+1,0)}={\bm{y}}_{j}$, with $j\sim\mathcal{U}(\llbracket 1,N\rrbracket)$ $\bm{y}_{i}^{(n+1,\,0)}={\bm{y}}_{i}$ $\bm{y}_{i}^{(n+1,\,0)}={\bm{y}}_{i}$
Random $\bm{y}^{(n+1,0)}\sim\mathcal{U}(\mathcal{Y})$ $\bm{y}_{i}^{(n+1,\,0)}\sim\mathcal{U}(\mathcal{Y})$ $\bm{y}_{i}^{(n+1,\,0)}\sim\mathcal{U}(\mathcal{Y})$
@@
B.4 Proposal distribution design for the DVRPTW
@@
The selected moves, designed for Local Search algorithms on vehicle routing problems (specifically for the PC-VRPTW for serve request and remove request), are given in Table˜5.
@@
In order to transform these deterministic moves into proposals, we first adapt the choice of clients $i$ and $j$, by sampling $i$ uniformly from $V^{1}_{s}(\bm{y})$, which contains the set of valid choices of client $i$ for move $s$ from solution $\bm{y}$. Then, we sample $j$ from $V^{2}_{s}(\bm{y})[i]\setminus\{i\}$ using the following softmax distribution: $P_{s}(j\mid i)=\frac{\exp\left[-d(i,j)/\beta\right]}{\sum_{k\in V^{2}_{s}(\bm{y})[i]\setminus\{i\}}\exp\left[-d(i,k)/\beta\right]}$, where $\beta>0$ is a neighborhood sampling temperature. The set $V^{2}_{s}(\bm{y})[i]$ contains all valid choices of client $j$ for move $s$ from solution $\bm{y}$, and is precised along with $V^{1}_{s}(\bm{y})$ in Table˜6. We normalize the distance measures inside the softmax, by dividing them by the maximum distance: $d(i,\cdot)\leftarrow d(i,\cdot)/\max_{k\in V^{2}_{s}(\bm{y})[i]\setminus\{i\}}d(i,k)$.
@@
relocate pair $\mathcal{D}(\bm{y})\setminus\bigl\{\mathcal{D}_{2}(\bm{y})\cup\mathcal{D}^{\text{last}}(\bm{y})\bigr\}$ $\mathcal{D}(\bm{y})\setminus\{\mathsf{next}(i)\}$
@@
serve request $\bar{\mathcal{D}}(\bm{y})$ $\mathcal{D}(\bm{y})\cup\mathcal{I}_{D}(\bm{y})$
@@
Table 6: Sets of valid clients for each move. $\mathcal{D}(\bm{y})$ contains all dispatched clients in solution $\bm{y}$. $\mathcal{D}_{1}(\bm{y})$ contains all dispatched clients that are the only client in their route. $\mathcal{D}_{2}(\bm{y})$ contains all dispatched clients that are in a route with 22 clients or less. $\mathcal{D}^{\text{last}}(\bm{y})$ contains all dispatched clients that are the last of their route. $\bar{\mathcal{D}}(\bm{y})$ contains all non-dispatched clients. $\mathcal{I}_{D}(\bm{y})$ contains the depot of the first empty route, if it exists (all routes may be non-empty), or else is the empty set. $\mathcal{I}_{1}(\bm{y})$ contains the only client in the last non-empty route if it contains exactly one client, or else is the empty set.
@@
  • For swap pair, the same arguments hold (leading to the same form for $q_{s}$), except for the equalities $V^{2}_{s}(\bm{y})[i]=V^{2}_{s}(\bm{y}^{\prime})[i]$ and $V^{2}_{s}(\bm{y})[j]=V^{2}_{s}(\bm{y}^{\prime})[j]$. Thus, we have the following form for the correction ratio:
    $\frac{q_{s}({\bm{y}}^{\prime},{\bm{y}})}{q_{s}({\bm{y}},{\bm{y}}^{\prime})}=\frac{\sum_{k\in V^{2}_{s}(\bm{y})[i]\setminus\{i\}}\exp\left[-d(i,k)/\beta\right]+\sum_{k\in V^{2}_{s}(\bm{y})[j]\setminus\{j\}}\exp\left[-d(j,k)/\beta\right]}{\sum_{k\in V^{2}_{s}(\bm{y}^{\prime})[i]\setminus\{i\}}\exp\left[-d(i,k)/\beta\right]+\sum_{k\in V^{2}_{s}(\bm{y}^{\prime})[j]\setminus\{j\}}\exp\left[-d(j,k)/\beta\right]}.$
@@
    $q_{s}(\bm{y},\bm{y}^{\prime})$ $=\frac{|\bigl\{\mathcal{D}(\bm{y})\setminus\mathcal{D}_{1}(\bm{y})\bigr\}\cup\mathcal{I}_{1}(\bm{y})|}{|\bigl\{\mathcal{D}(\bm{y})\setminus\mathcal{D}_{1}(\bm{y})\bigr\}\cup\mathcal{I}_{1}(\bm{y})|+|\bar{\mathcal{D}}(\bm{y})|}\times\frac{1}{|\bigl\{\mathcal{D}(\bm{y})\setminus\mathcal{D}_{1}(\bm{y})\bigr\}\cup\mathcal{I}_{1}(\bm{y})|}$
    $=\frac{1}{|\bigl\{\mathcal{D}(\bm{y})\setminus\mathcal{D}_{1}(\bm{y})\bigr\}\cup\mathcal{I}_{1}(\bm{y})|+|\bar{\mathcal{D}}(\bm{y})|}$
@@
    $q_{s}(\bm{y}^{\prime},\bm{y})$ $=\frac{1}{|\bigl\{\mathcal{D}(\bm{y}^{\prime})\setminus\mathcal{D}_{1}(\bm{y}^{\prime})\bigr\}\cup\mathcal{I}_{1}(\bm{y}^{\prime})|+|\bar{\mathcal{D}}(\bm{y}^{\prime})|}$
    $\times\frac{\exp\left[-\bar{d}(i)/\beta\right]}{\exp\left[-\bar{d}(i)/\beta\right]+\sum_{k\in\mathcal{D}(\bm{y}^{\prime})}\exp\left[-d(i,k)/\beta\right]}$
    $=\frac{1}{|\bigl\{\mathcal{D}(\bm{y})\setminus\mathcal{D}_{1}(\bm{y})\bigr\}\cup\mathcal{I}_{1}(\bm{y})|+|\bar{\mathcal{D}}(\bm{y})|}$
    $\times\frac{\exp\left[-\bar{d}(i)/\beta\right]}{\exp\left[-\bar{d}(i)/\beta\right]+\sum_{\begin{subarray}{c}k\in\mathcal{D}(\bm{y})\\ k\neq i\end{subarray}}\exp\left[-d(i,k)/\beta\right]}.$
@@
    $q_{s}(\bm{y}^{\prime},\bm{y})$ $=\frac{1}{|\bigl\{\mathcal{D}(\bm{y}^{\prime})\setminus\mathcal{D}_{1}(\bm{y}^{\prime})\bigr\}\cup\mathcal{I}_{1}(\bm{y}^{\prime})|+|\bar{\mathcal{D}}(\bm{y}^{\prime})|}$
    $\times\frac{\frac{1}{2}\cdot\exp\left[-d(i,\mathsf{prev}(i))\right]+\frac{1}{2}\cdot\exp\left[-d(i,\mathsf{next}(i))\right]}{\mathbf{1}_{\{\mathcal{I}_{D}(\bm{y}^{\prime})\neq\emptyset\}}\cdot\exp\left[-\bar{d}(i)/\beta\right]+\sum_{k\in\mathcal{D}(\bm{y}^{\prime})}\exp\left[-d(i,k)/\beta\right]}$
    $=\frac{1}{|\bigl\{\mathcal{D}(\bm{y})\setminus\mathcal{D}_{1}(\bm{y})\bigr\}\cup\mathcal{I}_{1}(\bm{y})|+|\bar{\mathcal{D}}(\bm{y})|}$
    $\times\frac{\frac{1}{2}\cdot\exp\left[-d(i,\mathsf{prev}(i))\right]+\frac{1}{2}\cdot\exp\left[-d(i,\mathsf{next}(i))\right]}{\mathbf{1}_{\{\mathcal{I}_{D}(\bm{y}^{\prime})\neq\emptyset\}}\cdot\exp\left[-\bar{d}(i)/\beta\right]+\sum_{\begin{subarray}{c}k\in\mathcal{D}(\bm{y})\\ k\neq i\end{subarray}}\exp\left[-d(i,k)/\beta\right]}.$
@@
    $q_{s}(\bm{y},\bm{y}^{\prime})$ $=\frac{|\bar{\mathcal{D}}(\bm{y})|}{|\bigl\{\mathcal{D}(\bm{y})\setminus\mathcal{D}_{1}(\bm{y})\bigr\}\cup\mathcal{I}_{1}(\bm{y})|+|\bar{\mathcal{D}}(\bm{y})|}$
    $\times\frac{\frac{1}{2}\cdot\exp\left[-d(i,j)\right]+\frac{1}{2}\cdot\exp\left[-d(i,j^{\prime})\right]}{\mathbf{1}_{\{\mathcal{I}_{D}(\bm{y})\neq\emptyset\}}\cdot\exp\left[-\bar{d}(i)/\beta\right]+\sum_{k\in\mathcal{D}(\bm{y})}\exp\left[-d(i,k)/\beta\right]}$
@@
    $q_{s}(\bm{y},\bm{y}^{\prime})$ $=\frac{1}{|\bigl\{\mathcal{D}(\bm{y})\setminus\mathcal{D}_{1}(\bm{y})\bigr\}\cup\mathcal{I}_{1}(\bm{y})|+|\bar{\mathcal{D}}(\bm{y})|}$
    $\times\frac{\exp\left[-\bar{d}(i)/\beta\right]}{\exp\left[-\bar{d}(i)/\beta\right]+\sum_{k\in\mathcal{D}(\bm{y})}\exp\left[-d(i,k)/\beta\right]}$
@@
    $q_{s}(\bm{y}^{\prime},\bm{y})=\frac{1}{|\bigl\{\mathcal{D}(\bm{y})\setminus\mathcal{D}_{1}(\bm{y})\bigr\}\cup\mathcal{I}_{1}(\bm{y})|+|\bar{\mathcal{D}}(\bm{y})|}.$
@@
In each case, we set $d(i,D)=\infty$ to account for the fact that the depot can never be sampled during the process (except in the serve request / remove request move, where we allow the depot of the first empty route / last non-empty route to be selected, for which we use the average distance to other requests as explained earlier) – in fact, the distance measure from a client to the depot is not even defined in the original HGS implementation.